46. The 2020, 2021, and 2022 partial balance sheets for Ottoman Manufacturing Company appear below. 2020 2021 2022 Current asNet credit sales for Ottoman were $126,000 for 2020, $120,000 for 2021 and $114,000 for 2022, while cost of goods sold was $849. The 2021 and 2022 balance sheets for Ottoman Manufacturing Company appear below: 2021 2022 Current Assets Cash Accounts rCurrent Assets Cash Accounts receivable, net Inventory $ 16,000 $ 15,000 20,000 19.000 30.000 28.000 66,000 62,000 60,000 40.
Answer #1

Q 46: Account Receivable Turnover = Net cr. Sales/average receivable= 120000/24000= 5

Q 47: Inventory Turnover=COGS/Avg. Inventory= 82400/26000=3.17

Q 48: days to sell inventory = 365days/Inventory turnover ratio = 365/2.5= 146 days

Q 49: Debt Ratio =Total liabilities/Total Assets = 70000/102000 = 68.6%